# Dynamical thermalization, Rayleigh-Jeans condensate, vortexes and wave collapse in quantum chaos fibers and fluid of light

### v2

@@ V.1 Thermalization features
In Figure 13 we show the probability distribution $<\rho_{m}(t)>$ over linear eigenmodes $\psi_{m}$ averaged over a long time interval $500\leq t\leq 1000$ for the same initial states as in Figure 12 at $\beta=10$. For all initial states we have a high fraction of RJ condensate at the linear ground state $\psi_{g}=\psi_{1}$. The probabilities $\rho_{m}$ at higher modes with $m>1$ approximately follow the RJ thermal distribution (2) which is characterized by a decay $\rho_{m}\sim 1/(E_{m}-\mu)$ shown by a dashed black curve for initial energy $E=20$. The BE ansatz at the same energy $E=20$ is also shown in Figure 12 by the dashed red curve and it is obviously very different from the presented numerical results. Even if there are visible fluctuations of $\rho_{m}$ values, which we attribute to not very long times reached in heavy numerical simulations, on average we show that the dynamical thermalization and chaos in the defocucing fiber NSE (3) leads to the RJ thermal distribution (2) providing a good description of numerical results including the formation of RJ condensate.
@@
The Boltzmann H-theorem of 1872 boltzmann1 , see also e.g. mayer ; landau , states that during a thermalization process a system entropy $S(t)$ cannot decrease with time $t$, thus $S(t)$ can only grow with time or stay constant at its maximal value in a thermal equilibrium. In our system (3) with the dynamical thermalization we have a strikingly different situation. The results presented in Figure 14 obtained at the same parameters as in Figures 1213 show that the system entropy $S(t)=-\sum_{m}\rho_{m}(t)\ln\rho_{m}(t)$ initially grows with time, reaches a maximal value and then decreases up to a certain smaller value (being up to a factor 3 smaller than at maximum) corresponding to the RJ thermal distribution (2) with RJ condensate. Our qualitative explanation of this surprising behavior of $S(t)$ is the following: initially $S(t\approx 0)\approx\ln 2$ since initially two linear eigenmodes are populated, then at higher times transitions to other linear eigenmodes are produced by nonlinear $\beta$ term and $S$ is increased reaching its maximal value, but then a slow thermalizarion due to chaos leads to the RJ thermal distribution (2) where the fraction of RJ condensate is close to unity (see Figure 12), since RJ condensate fraction is so high the entropy $S$ of steady-state has a relatively small value being significantly smaller than its maximal value. We argue that there is no contradiction with the Boltzmann H-theorem since it is stated for systems with only one energy integral of motion, while in our system we have two integrals of energy and norm.
@@
Figure 14: Dependence of entropy $S$ on time $t$ for $\beta=10$ and other parameters as in Figure 12; values of $S$ are averaged in the time window $\delta t=5$ to reduce the fluctuations; short time behavior of $S(t)$ is shown in Appendix Figure A2.
@@
Figure 15: Entropy $S$ vs energy $E$ are shown by points for the initial states of Figure 12 and Figure 13. The left panel displays the time average for times between 30 and 50, while the right panel shows the average for times between 500 and 1000; here as in Figure 12 $E$ is the total energy. In both panels, the curves are shown for the Bose-Einstein (BE) and Rayleigh-Jeans (RJ) distributions for the billiard with $N=1000$.
@@
In Figure 15 we compare the dependence of entropy $S$ values, averaged over a certain time interval at low and high times, on initial energy of linear modes $E_{m}$. On short times $30\leq t\leq 40$ this dependence $S(E_{m})$ is close to those of one given by the BE thermal distribution (1) while at longer times $500\leq t\leq 1000$ it is more close to those of RJ thermal distribution (2). The reason of this is related to the unusual entropy time dependence discussed for Figure 14 where $S$ grows at initial times and later decays to a smaller value induced by the RJ thermal condensate. On the basis of obtained results we argue that the result of eplbill ; sinaiosc that the dependence $S(E)$ is close to those of the BE distribution (1) instead of RJ one (2) is related to significantly shorter times reached there. Here the FREEFEM codes allow us to reach times being by a factor 30 longer where the effects of dynamical RJ thermalization and RJ condensate start to play the dominant role.
@@ V.2 Stability island near the ground state linear eigenmode
Figure 17: Stability of self-consistent states obtained near the ground state and first excited state: the error of initial deviation remains approximately constant in time $t$ showing that the state $\psi_{s}$ is table near a vicinity of the ground state while in a vicinity of the first excited state $\psi_{s2}$ the error grows exponentially with time showing that the corresponding self-consistent state is unstable.
@@ V.3 Quasi-integrable KAM regime
According to the KAM theorem arnold ; sinai ; chirikov1979 ; lichtenberg ; kuksin we expect that at small nonlinearity $\beta\ll\beta_{ch}$ the NSE dynamics is integrable in the major part of system phase space, even if certain small chaotic regions can remain, e.g. around thin separatrix layers with the Arnold diffusion. Indeed, the results of Figure 18 show that at small $\beta$ the entropy $S(t)$ is not growing with time showing only regular oscillations at $\beta=1$ or almost constant small value $S<1$ at $\beta=0.5$. Even if $S$ values at $\beta=1;0.5$ are comparable with $S$ value at long time $t=1000$ for $\beta=10$ the situation is qualitatively different: at $\beta=10$ the dynamics is chaotic leading to RJ thermalization with a high fraction of RJ condensate (that gives small $S$ values at $t=1000$) while at $\beta\leq 1$ the probabilities $\rho_{m}(t)$ are simply regularly oscillating near there initial values as it is shown in Appendix Figure  A3.
@@
Figure 18: Entropy $S$ dependence on time for the initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$ and $\beta=0.5;1;10$ (values of $S$ are averaged over a time window $\delta t=5$).
@@ VI Wave collapse for focusing fiber NSE
In the pre-collapse regime at $C=0.18$ there is an interesting behavior of wave function $\psi(x,y,t)$ that can be seen by its projection probability ${\rho^{*}}_{n}$ on the eigenbasis $1\leq n<\infty$ of the Schrödinger equation (3) in which the nonlinear term is considered as the instantaneous effective attractive potential $U_{eff}=\beta|\psi(x,y,t)|^{2}\leq 0$. The dependence of ${\rho^{*}}_{n}$ on time $t$ is shown in the bottom panel of Figure 20. At small times $t<0.1$ there is about ${\rho^{*}}_{1}=0.95$ to $0.915$ probability located at the ground state $n=1$ of the instantaneous potential $U_{eff}(x,y,t)$. But $U_{eff}$ is changing with time that leads to a transition of almost all fraction of ${\rho^{*}}_{1}$ to other eigenstate at $n=2$ at higher moment of time $t$. Then with growth of $t$ there are transitions to high and higher states $n$ of instantaneous potential $U_{eff}$. The maximal $n$ values (up to $n=16$) are reached at the minimal size of wave function corresponding to the maximal values of $E_{kin}$ in the top panel. Then, when the packet squeezing passed, its maximal value and the packet size starts to increase the population ${\rho^{*}}_{n}$ starts again go to minimal $n$ values returning back to almost all probability at the ground state $n=1$ of $U_{eff}$ when the packet returns of its maximal site (minimal of $E_{kin}$ in the top panel). It is interesting to note that the transitions between states $n$ happens in diabatic manner with almost all probability transferred from one $n$ value to another one. These results show that in the pre-collapse regime there is rather nontrivial time evolution of wave function with its oscillating size. It would be interesting to study this behavior in more detail but this goes outside of scopes of this work.
@@ VII Dynamical thermalization in focusing fiber NSE
Figure 21: Dependence of ground state probability $\rho_{g}(t)$ on time $t$ as in Figure 12 but for $\beta=-10$ and only one initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$; the projection probability on the linear ground state $\psi_{1}$ is shown by blue curve and those on the ground state of instantaneous potential $U_{eff}=\beta|\psi(t)|^{2}$ is shown by green curve; total energy is $E=60.36$; the dashed lines show BE ansatz (red) and RJ ansatz (black) with the same values of $E$, $T$, $\mu$, $N$ as in Figure 12.
@@
Figure 22: Projected probabilities $<\rho_{m}>$ in the linear eigenstates $\psi_{m}$ and averaged in the time interval $400\leq t\leq 800$ are shown in dependence on linear eigenenergies $E_{m}$ at $\beta=-10$ and initial state as in Figure 21; dashed curves show the BE and RJ ansatz with same parameters as in Figure 12.
@@
Figure 23: Dependence of entropy $S(t)=-\sum_{m}\rho_{m}(t)\ln\rho_{m}(t)$ on time $t$ for $\beta=-10$ (blue curve ) and $\beta=-0.5$ (red curve) with initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$; left/right panels show long/short time behavior.
The time dependence of entropy $S(t)$ at $\beta=-10;-0.5$ and initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$ is shown in Figure 23. This dependence is similar to those in Figure 14 at $\beta=10$. At small $\beta=-0.5$ the entropy remains practically unchanged corresponding to the KAM regime. In this quasi-integrable KAM regime at $\beta=-0.5$ only regular probability oscillations in time (see Appendix Figure A6).
@@ IX NSE dynamics at long times
[…]
@@ X Experimental parameters for a quantum chaos fiber
Here we discuss the requested experimental parameters for D-shape fiber to have sufficiently strong nonlinearity $\beta$ to reach regime of dynamical thermalization. We start from experimental conditions already realized in aplbabin ; wabnitz with a circle fiber cross-section where the classical dynamics of linear system at $\beta=0$ in (3) is integrable. These fiber experiments were done at high laser beam intensity and the theoretical description of the complex field envelope $A(x,y,z,t)$ expressed in $[\sqrt{W}/m]$ is described by the NSE type equation in which we keep notations of wabnitz :
@@
that gives $\beta\approx 0.32$ for typical parameters of wabnitz with $I=10GW/cm^{2}$, $n_{2}=\times 10^{-20}m^{2}/W$, $n_{co}=1.5$, $\lambda=1000nm$, $r_{c}=30\mu m$. In (20) the length unit is $z_{0}={r_{c}}^{2}/\lambda\approx 0.1cm$ for the above values $r_{c}$, $\lambda$. Thus for a fiber length of $1$ meter we have the dimensionless length of $z/z_{0}\approx 1000$ that corresponds to the dimensional time interval $t$ of NSE (3) being also 1000. In our numerical simulations of NSE (3) we are reaching similar time scales $t\approx 100$ that is comparable with the fiber length of $0.4m$ in wabnitz .
@@
XI Discussion and conclusion
Our numerical and analytical studies of the NSE fiber light fluid dynamics in a chaotic D-shape billiard show that it is characterized by various unusual and rich phenomena. Above the chaos border at $|\beta|>\beta_{ch}\sim 10$ the dynamical thermalization leads to the RJ thermal distribution (2) with a formation of the RJ condensate in the ground state of linear system which fraction has about 80-90% of total probability. Certain similarities between RJ and Froöhlich condensates are discussed. We show that at $\beta>0$ near the ground state there is a self-consistent state with a certain island of stable motion around it. In contrast to the Boltzmann H-theorem boltzmann1 ; mayer ; landau in the chaotic regime the entropy time evolution $S(t)$ is characterized by initial increase at short times and later relaxation from its maximal value to a significantly smaller value on long times in the steady-state regime that is explained by the formation of RJ condensate. We point that the presence of two integrals of motion, energy and norm, is at the origin of such a difference from the H-theorem formulated for systems with only one energy integral of motion. In absence of nonlinearity our system belongs to the systems of quantum chaos, thus having many features as in the Random Matrix Theory. On these grounds we argue that the dynamical thermalization induced by nonlinearity represents a generic features of such thermalization similar to those discussed for nonlinear perturbation of Random Matrix theory in rmtprl . We also note that even if the phenomenon of RJ condensation in nonlinear systems has been discussed for multi-mode fibers picozzi1 ; picozzi2 ; picozzi3 ; picozzi2024 its origin was not associated with the emergence of dynamical chaos above a certain chaos border below which the RJ thermalization and RJ condensate are absent and dynamics is quasi-integrable as in the KAM theory. We attribute the so-called self-cleaning phenomenon observed in multi-mode optical fibers aplbabin ; wabnitz to the formation of RJ condensate induced by dynamical thermalization and chaos. A pictorial image of time development of dynamical thermalization and RJ condensate formation is shown in Figure 27.

### v3

@@ V.1 Thermalization features
In Figure 13 we show the probability distribution $\rho_{m}(t)>$ over linear eigenmodes $\psi_{m}$ averaged over a long time interval $500\leq t\leq 1000$ for the same initial states as in Figure 12 at $\beta=10$. For all initial states we have a high fraction of RJ condensate at the linear ground state $\psi_{g}=\psi_{1}$. The probabilities $\rho_{m}$ at higher modes with $m>1$ approximately follow the RJ thermal distribution (2) which is characterized by a decay $\rho_{m}\sim 1/(E_{m}-\mu)$ shown by a dashed black curve for initial energy $E=20$. The BE ansatz at the same energy $E=20$ is also shown in Figure 12 by the dashed red curve and it is obviously very different from the presented numerical results. Even if there are visible fluctuations of $\rho_{m}$ values, which we attribute to not very long times reached in heavy numerical simulations, on average we show that the dynamical thermalization and chaos in the defocusing fiber NSE (3) leads to the RJ thermal distribution (2) providing a good description of numerical results including the formation of RJ condensate.
@@
The Boltzmann H-theorem of 1872 boltzmann1 , see also e.g. mayer ; landau , states that during a thermalization process in a classical system entropy $S_{B}(t)$ cannot decrease with time $t$, thus $S_{B}(t)$ can only grow with time or stay constant at its maximal value in a thermal equilibrium. For the quantum von Neumann entropy $S_{q}(t)$ we have a different situation in our system (3) with the dynamical thermalization and RJ condensation. Thus the results presented in Figure 14, obtained at the same parameters as in Figures 1213, show that the quantum von Neuman entropy $S_{q}(t)=-\sum_{m}\rho_{m}(t)\ln\rho_{m}(t)$ initially grows with time, reaches a maximal value and then decreases up to a certain smaller value (being up to a factor 3 smaller than at maximum) corresponding to the RJ thermal distribution (2) with RJ condensate.
Our qualitative explanation of this surprising behavior of $S_{q}(t)$ is the following: initially $S_{q}(t\approx 0)\approx\ln 2$ since initially two linear eigenmodes are populated, then at higher times transitions to other linear eigenmodes are produced by nonlinear $\beta$ term and $S_{q}$ is increased reaching its maximal value, but then a slow thermalization due to chaos leads to the RJ thermal distribution (2) where the fraction of RJ condensate is close to unity (see Figure 12), since RJ condensate fraction is so high the entropy $S_{q}$ of steady-state has a relatively small value being significantly smaller than its maximal value.
@@
Figure 14: Dependence of quantum von Neumann entropy $S_{q}(t)=-\sum_{m}\rho_{m}(t)\ln\rho_{,}(t)$ on time $t$ for $\beta=10$ and other parameters as in Figure 12; values of $S_{q}$ are averaged in the time window $\delta t=5$ to reduce the fluctuations; short time behavior of $S_{q}(t)$ is shown in Appendix Figure A2.
The above time dependence of $S_{q}(t)$ is not in a contradiction with the Boltzmann H-theorem formulated for the Boltzmann entropy $S_{B}(t)=\sum_{m}\ln\rho_{m}(t)$. However, the time dependence of the Boltzmann entropy $S_{B}(t)$ shown in Appendix Figure A3 is also non-monotonic in contradiction with the H-theorem. The maximal value of $S_{B}$ is about 50% higher than its value at high times where the thermalization takes place. Such a behavior is found bith for $\beta>0$ and $\beta<0$ (see Appendic Figure A3). At the same time we should point on the problem of the definition of $S_{B}$ in our system (3) with infinite number of linear eigenmodes $1\leq m\leq\infty$. Indeed, it is clear that probabilities $\rho_{m}$ at high modes should be small that leads to a formal divergence for $S_{B}$ value given by the above expression. Due to this reason we present $S_{B}(t)$ time dependence in Appendix Figure A3 where $S_{B}$ is computed by a summation over a finite interval of eigenmodes with $1\leq m\leq m_{max}$. The change of $m_{max}$ does not eliminate the non-monotonic time dependence of $S_{B}(t)$. Strongly negative values of $S_{B}$ can be in principle compensated by an arbitrary constant $C$ in the definition of $S_{B}$ in 7). There is also the problem of $S_{B}$ definition at small initial times $t\ll 1$ but it can be avoided by taking average over a time interval that in any case should be used to suppress fluctuations.
We note that the numerical simulations of defocusing NSE presented in picozzi3 show a monotonic increase of $S_{B}(t)$. However, the numerical simulations there were done with a finite number of about 40 modes present in a fiber. Our numerical simulations here with FREEFEM codes have much larger number of eigenmodes that may be at the origin if non-monotonic dependence of $S_{B}(t)$ found here in Appendix Figure A3. Formally, it is possible to say that the H-theorem is formulated for a gas of many particles and there is no guaranty that it should be valid for the NSE case of our system (3). At the same time we should note that our system (3) in the thermalized regime is always located in the RJ condensate phase since number of levels is formally unlimited here. It is also possible that we do not reach complete thermalization on very high eigenmodes $m$ that gives a significant contribution to $S_{B}$ that may be also responsible for non-monotonic time dependence of $S_{B}(t)$. We restrict ourselves by pointing on the non-monotonic time dependence of Boltzmann entropy $S_{B}(t)$ and leaving the resolution of this problem to future studies.
@@
Figure 15: Dependence of quantum von Neumann entropy $S_{q}$ on energy $E$ is shown by points for the initial states of Figure 12 and Figure 13. The left panel displays the time average for times between 30 and 50, while the right panel shows the average for times between 500 and 1000; here as in Figure 12 $E$ is the total energy. In both panels, the curves are shown for the Bose-Einstein (BE) and Rayleigh-Jeans (RJ) distributions for the billiard with $N=1000$.
@@
In Figure 15 we compare the dependence of quantum von Neumann entropy $S_{q}$ values, averaged over a certain time interval at low and high times, on initial energy of linear modes $E_{m}$. On short times $30\leq t\leq 40$ this dependence $S_{q}(E_{m})$ is close to those of one given by the BE thermal distribution (1) while at longer times $500\leq t\leq 1000$ it is more close to those of RJ thermal distribution (2). The reason of this is related to the unusual entropy time dependence discussed for Figure 14 where $S$ grows at initial times and later decays to a smaller value induced by the RJ thermal condensate. On the basis of obtained results we argue that the result of eplbill ; sinaiosc that the dependence $S_{q}(E)$ is close to those of the BE distribution (1) instead of RJ one (2) is related to significantly shorter times reached there. Here the FREEFEM codes allow us to reach times being by a factor 30 longer where the effects of dynamical RJ thermalization and RJ condensate start to play the dominant role.
@@ V.2 Stability island near the ground state linear eigenmode
Figure 17: Stability of self-consistent states obtained near the ground state and first excited state: the error of initial deviation remains approximately constant in time $t$ showing that the state $\psi_{s}$ is taken near an integrable vicinity of the ground state while in a vicinity of the first excited state $\psi_{s2}$ the error grows exponentially with time showing that the corresponding self-consistent state is unstable.
@@ V.3 Quasi-integrable KAM regime
According to the KAM theorem arnold ; sinai ; chirikov1979 ; lichtenberg ; kuksin we expect that at small nonlinearity $\beta\ll\beta_{ch}$ the NSE dynamics is integrable in the major part of system phase space, even if certain small chaotic regions can remain, e.g. around thin separatrix layers with the Arnold diffusion. Indeed, the results of Figure 18 show that at small $\beta$ the entropy $S(t)$ is not growing with time showing only regular oscillations at $\beta=1$ or almost constant small value $S<1$ at $\beta=0.5$. Even if $S$ values at $\beta=1;0.5$ are comparable with $S$ value at long time $t=1000$ for $\beta=10$ the situation is qualitatively different: at $\beta=10$ the dynamics is chaotic leading to RJ thermalization with a high fraction of RJ condensate (that gives small $S$ values at $t=1000$) while at $\beta\leq 1$ the probabilities $\rho_{m}(t)$ are simply regularly oscillating near there initial values as it is shown in Appendix Figure  A4.
We use here he term quasi-integrable KAM regime to point that even in the limit of very small nonlinearity there are small regions of chaotic component, usually around separatrix curves (see e.g. chirikov1979 ; lichtenberg ).
@@
Figure 18: Time dependence of the quantum von Neumann entropy $S_{q}(t)$ for the initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$ and $\beta=0.5;1;10$ (values of $S$ are averaged over a time window $\delta t=5$).
@@ VI Wave collapse for focusing fiber NSE
In the pre-collapse regime at $C=0.18$ there is an interesting behavior of wave function $\psi(x,y,t)$ that can be seen by its projection probability ${\rho^{*}}_{n}$ on the eigenbasis $1\leq n<\infty$ of the Schrödinger equation (3) in which the nonlinear term is considered as the instantaneous effective attractive potential $U_{eff}=\beta|\psi(x,y,t)|^{2}\leq 0$. The dependence of ${\rho^{*}}_{n}$ on time $t$ is shown in the bottom panel of Figure 20. At small times $t<0.1$ there is about ${\rho^{*}}_{1}=0.95$ to $0.915$ probability located at the ground state $n=1$ of the instantaneous potential $U_{eff}(x,y,t)$. But $U_{eff}$ is changing with time that leads to a transition of almost all fraction of ${\rho^{*}}_{1}$ to other eigenstate at $n=2$ at higher moment of time $t$. Then with growth of $t$ there are transitions to high and higher states $n$ of instantaneous potential $U_{eff}$. The maximal $n$ values (up to $n=16$) are reached at the minimal size of wave function corresponding to the maximal values of $E_{kin}$ in the top panel. Then, when the packet squeezing passed, its maximal value and the packet size starts to increase the population ${\rho^{*}}_{n}$ starts again go to minimal $n$ values returning back to almost all probability at the ground state $n=1$ of $U_{eff}$ when the packet returns of its maximal site (minimal of $E_{kin}$ in the top panel). It is interesting to note that the transitions between states $n$ happens in adiabatic manner with almost all probability transferred from one $n$ value to another one. These results show that in the pre-collapse regime there is rather nontrivial time evolution of wave function with its oscillating size. It would be interesting to study this behavior in more detail but this goes outside of scopes of this work.
@@ VII Dynamical thermalization in focusing fiber NSE
Figure 21: Dependence of ground state probability $\rho_{g}(t)$ on time $t$ as in Figure 12 but for $\beta=-10$ and only one initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$; the projection probability on the linear ground state $\psi_{1}$ is shown by blue curve and those on the ground state of instantaneous potential $U_{eff}=\beta|\psi(t)|^{2}$ is shown by green curve; total energy is $E=60.36$; the dashed lines show BE ansatz (red) and RJ ansatz (black) with the same values of $E$, $T$, $\mu$, NN as in Figure 12.
@@
Figure 22: Projected probabilities $\rho_{m}>$ in the linear eigenstates $\psi_{m}$ and averaged in the time interval $400\leq t\leq 800$ are shown in dependence on linear eigenenergies $E_{m}$ at $\beta=-10$ and initial state as in Figure 21; dashed curves show the BE and RJ ansatz with same parameters as in Figure 12.
@@
Figure 23: Dependence of the quantum von Neumann entropy $S_{q}(t)=-\sum_{m}\rho_{m}(t)\ln\rho_{m}(t)$ on time $t$ for $\beta=-10$ (blue curve ) and $\beta=-0.5$ (red curve) with initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$; left/right panels show long/short time behavior. The time dependence of classical Boltzamann entropy $S_{B}(t)=\sum_{m}\ln\rho_{m}(t)$ is shown in Appendix Figure A3 at the same system parameters.
The time dependence of quantum von Neumann entropy $S_{q}(t)$ at $\beta=-10;-0.5$ and initial state $\psi(t=0)=(\psi_{m}+\psi_{m+1})/\sqrt{2}$ at $m=11$ is shown in Figure 23. This dependence is similar to those in Figure 14 at $\beta=10$. At small $\beta=-0.5$ the entropy $S_{q}$ remains practically unchanged corresponding to the KAM regime. In this quasi-integrable KAM regime at $\beta=-0.5$ only regular probability oscillations in time (see Appendix Figure A7).
@@ IX NSE dynamics at long times
Here $C_{m}(t)$ are wave function amplitudes in the eigenbasis $\psi_{m}$ so that $\psi(x,y,t)=\sum_{m}C_{m}(t)\psi_{m}(x,y)$ and interaction matrix elements induced by nonlinearity are
$V_{{m}{m_{1}}{m_{2}}{m_{3}}}=\int_{x,y}{\psi^{*}}_{m}\psi_{m_{1}}{\psi^{*}}_{m_{2}}\psi_{m_{3}}dxdy\;.$ (23)
@@ X Experimental parameters for a quantum chaos fiber
Here we discuss the requested experimental parameters for D-shape fiber to have sufficiently strong nonlinearity $\beta$ to reach regime of dynamical thermalization. We start from experimental conditions already realized in aplbabin ; wabnitz with a circle fiber cross-section where the classical dynamics of linear system at $\beta=0$ in (3) is integrable. These fiber experiments were done at high laser beam intensity and the theoretical description of the complex field envelope $A(x,y,z,t)$ expressed in $\sqrt{W}/m]$ is described by the NSE type equation in which we keep notations of wabnitz :
@@
that gives $\beta\approx 0.32$ for typical parameters of wabnitz with $I=10GW/cm^{2}$, $n_{2}=3.2\times 10^{-20}m^{2}/W$, $n_{co}=1.5$, $\lambda=1000nm$, $r_{c}=30\mu m$. In (24) the length unit is $z_{0}={r_{c}}^{2}/\lambda\approx 0.1cm$ for the above values $r_{c}$, $\lambda$. Thus for a fiber length of $1$ meter we have the dimensionless length of $z/z_{0}\approx 1000$ that corresponds to the dimensional time interval $t$ of NSE (3) being also 1000. In our numerical simulations of NSE (3) we are reaching similar time scales $t\approx 100$ that is comparable with the fiber length of $0.4m$ in wabnitz .
@@
XI Discussion
Our numerical and analytical studies of the NSE fiber light fluid dynamics in a chaotic D-shape billiard show that it is characterized by various unusual and rich phenomena. Above the chaos border at $|\beta|>\beta_{ch}\sim 10$ the dynamical thermalization leads to the RJ thermal distribution (2) with a formation of the RJ condensate in the ground state of linear system which fraction has about 80-90% of total probability. Certain similarities between RJ and Froöhlich condensates are discussed. We show that at $\beta>0$ near the ground state there is a self-consistent state with a certain island of stable motion around it.
In the regime of chaotic dynamical thermalization with emergence of RJ condensate we show that the time dependences of the quantum von Neumann entropy $S_{q}(t)$ and the classical Bolzmann entropy $S_{B}(t)$ are non-monotonic in time being characterized by an initial increase at short times and later relaxation from its maximal value to a significantly smaller value on long times in the steady-state regime with RJ condensate. Such a result for $S_{B}(t)$ is in contrast to the Boltzmann H-theorem boltzmann1 ; mayer ; landau even if no such theorem has been proven for the NSE equation in a quantum chaos fiber. We point on a certain divergence problem of $S_{B}$ definition in our system (3) due to presence of unlimited number of billiard linear eigenmodes assuming that this can be at the origin of this contrast with the H-theorem.
In absence of nonlinearity our system belongs to the systems of quantum chaos, thus having many features as in the Random Matrix Theory. On these grounds we argue that the dynamical thermalization induced by nonlinearity represents a generic features of such thermalization similar to those discussed for nonlinear perturbation of Random Matrix theory in rmtprl . We also note that even if the phenomenon of RJ condensation in nonlinear systems has been discussed for multi-mode fibers picozzi1 ; picozzi2 ; picozzi3 ; picozzi2024 its origin was not associated with the emergence of dynamical chaos above a certain chaos border below which the RJ thermalization and RJ condensate are absent and dynamics is quasi-integrable as in the KAM theory. We attribute the so-called self-cleaning phenomenon observed in multi-mode optical fibers aplbabin ; wabnitz ; fiber1 to the formation of RJ condensate induced by dynamical thermalization and chaos. A pictorial image of time development of dynamical thermalization and RJ condensate formation is shown in Figure 27.